# Probing AGN Feedback in Dwarf Galaxies with Spatially Resolved NIR Coronal Lines from JWST

### v1

@@ 5.4 Outflows traced by Coronal Lines
Assuming the secondary component seen in the CLs is an outflow, we can derive standard estimates of energetics, including the mass outflow and kinetic energy rates. Previous studies on CL energetics (2011ApJ...739...69M; 2017MNRAS.470.2845R) account for the velocities and the extents of the outflow. However, they do not measure the mass from the luminosities of the CL, but instead assume a certain filling factor and make use of the hydrogen mass alone. The unknown filling factor leads to large uncertainties in the calculated energetics. In this work, we make use of [Si VI] luminosities in order to estimate the amount of gas mass present in the outflow, and use that to determine the outflow energetics (see Appendix A for a more detailed derivation). We use the following equation to derive the mass outflow rates for a steady state-outflow:
@@
where is the extinction-corrected [Si VI] luminosity and is the electron density, measured from the [S II] 6716, 6731 line ratios, obtained from Gemini/GMOS spectra or Keck/LRIS spectra (2020ApJ...905..166L). We caution that the we adopt from the [S II] doublet likely traces lower-ionization NLR gas rather than the coronal line–emitting region itself (2019MNRAS.486.4290B; 2020MNRAS.498.4150D). If the CLR is denser than the [S II]-emitting gas, our assumed is underestimated and the inferred outflow masses and rates should be treated as upper limits.
@@
Note. — Column (1): Short name of the target. Column (2): Individual outflow components from the best fits. Column (3) Electron density measured from the [S II] 6716, 6731 line ratio based on the total line flux from the spatially integrated, Gemini/GMOS spectra or Keck/LRIS spectra (see 2020ApJ...905..166L). Column (4): Outflow radius adopted in the calculation of mass, momentum, and kinetic energy outflow rates when the outflows are spatially resolved. Column (5): Average offset velocity calculated from the best fits. Column (6): Average W80 calculated from the best fits. Column (7): Ionized gas mass of the corresponding outflow component. Column (8): Ionized gas mass outflow rate of the corresponding outflow component. Column (9): Ionized gas kinetic energy outflow rate of the corresponding outflow component. Column (10): Ionized gas momentum outflow rate of the corresponding velocity component
The calculations of the mass, momentum, and kinetic energy outflow rates depend on the spatial extent of the outflows which we calculate in Section 3.4 and discuss in Section 5.2. Specifically, the energetics are calculated by summing up quantities over individual spaxels:
@@
In Figure 15, we plot the mass and kinetic energy outflow rates estimated for the outflowing gas traced by the CLs as a function of mass and AGN luminosity. We also compare the outflow rates from CL to the ionized [O III] gas measurements and find that the CL outflow rates are 1 dex lower, which could be due to underestimating the W80 values due to the de-wiggling process (see Section 3.3). We also compare the energetics measured in this work to energetics calculated using [Si VI] outflows from ground-based low-z Seyferts (2011ApJ...739...69M; 2017MNRAS.470.2845R; 2018MNRAS.481L.105M). Since the mass outflow rates in these works were calculated assuming a filling factor, we also calculated the energetics from the AGN dwarfs using the equations presented in 2011ApJ...739...69M (indicated by the boxes with a cross). We assume a filling factor (0.001) similar to 2011ApJ...739...69M and find that this measurement over estimates the values for energetics by at least two orders of magnitude. Even if we assume higher ne and filling factor values suitable for dwarf galaxies, the energetics are still found to be higher. This could indicate that the filling factors for CL are much lower than previously assumed.
Because the highly ionized phase of the gas traced by CLs appears neither fast nor radially extended, it likely cannot expel gas from the galaxy. However, its energetics may be sufficient to disturb gas in the central regions, potentially contributing to feedback-driven gas redistribution that has been invoked to produce cored (rather than cuspy) dark-matter profiles in dwarf galaxies (2010AdAst2010E...5D). They might also accelerate gas and redistribute mass in the central, high–gravitational-potential regions, potentially driving more extended, lower-ionization outflows (e.g., [O III]).
6 Summary and Conclusions
Using JWST/NIRSpec, we undertake the first spatially resolved examination of NIR coronal lines in dwarf galaxies with AGN and ionized-gas outflows to address the challenges in unambiguously identifying AGN activity and assessing AGN-driven feedback in these low-mass systems. We generate spatial maps of the flux and velocity of the different coronal lines, allowing us to examine in detail the influence of AGN feedback, as traced by the coronal lines, on the host dwarf galaxies. We present our main results as follows:
@@ 6 Summary and Conclusions
Based on shock and AGN models, we find that AGN photoionization is largely responsible for producing CLs. J1009, in particular, is dominated by AGN photoionization with minimal to no contribution from shock ionization to producing the CLs. The models also suggest that J1009 is hosting a highly accreting, low-mass black hole that is capable of producing numerous CL with I.P.s as high as 447 eV.
@@
The average extents of the CLs across the four dwarf galaxies are 0.5 kpc, which is similar to the extents found tracing NIR CL in massive galaxies. The small sizes of dwarf galaxies mean their CLRs can occupy 10% of the host galaxy, about ten times the fractional extent seen in massive galaxies. In J0954, an outflow radius of 0.5 kpc estimated from blueshifted UV absorption (2024ApJ...965..152L) agrees with the CL extents, supporting CLs as tracers of the ionized outflow.
@@
The presence, extent, and luminosity of the CLs correlate with the properties of the ionized gas outflows traced by [O III]. In galaxies with spatially resolved CL emission, CLs exhibit biconical structure, align with the outflow axis, consistent with an AGN-powered outflow. The CL luminosities correlate with both the ionized gas outflow luminosities and the FWHM of the outflows. Two of the galaxies with CL detections also have small dust extinction values, suggesting that AGN-driven outflows may enhance CL emission by clearing dust from the nuclear regions.
@@
Several CLs exhibit wings, suggestive of a possible secondary kinematic component. If this is an outflow, energy estimates imply this highly ionized gas can have a dynamical impact on the gas near galaxy centers.
@@
The results confirm that the outflows detected in these dwarf galaxies are predominantly AGN-powered and that CLs may trace their presence and impact. By revealing signatures of dust clearing, gas heating, and kinematically distinct ionized outflows, CLs may provide a direct view of how AGN in dwarf galaxies can exert feedback, potentially influencing their evolution. Our findings highlight the potential for CLs as diagnostics of AGN activity in low-mass, dust-poor environments, which are conditions reminiscent of galaxies in the early universe. With increased observations of AGN in faint galaxies at high redshifts (2023MNRAS.525.2087B; 2023ApJ...952..142F; 2023ApJ...959...39H; 2023ApJ...954L...4K; 2023Natur.616..266L; 2024Natur.627...59M), including detections of CLs such as [Ne V] in dwarf galaxies at the epoch of reionization (2024MNRAS.534.2633C) using JWST, understanding the formation and detection of CLs in dwarf galaxies is critical. Given that local dwarf galaxies are crucial analogs for understanding the formation and evolution of the earliest galaxies in the high-redshift universe, where they are increasingly recognised as key drivers of reionization (2024Natur.626..975A) and where AGN feedback in dwarf galaxies is gaining importance in galaxy quenching (2024A&A...690A.286A), these results provide invaluable empirical constraints for cosmological simulations and future JWST observations aimed at unveiling the role of AGN in early galaxy evolution.
@@ Appendix A Outflow energetics
Spatially resolved IFU data has proven to be integral to deriving various energetics of outflows (the amount of mass carried by the outflows, the rates of mass, momentum and kinetic energy of the outflow), by providing constraints on the spatial extent of the outflow, which is a crucial factor in the calculation of the energetics. The energetics are dependent on several assumptions, such as the shape of the outflow (spherical vs. biconical). The ionized gas mass of the outflows can be calculated based on either the [O III] 4959, 5007 line luminosity or the Balmer line (H or H) luminosity of the outflowing, line emitting gas. Calculating the coronal gas mass using CLs is challenging. Previous works (2011ApJ...739...69M; 2017MNRAS.470.2845R; 2018MNRAS.481L.105M; 2020ApJ...895L...9R), typically assume a spherically symmetric outflow with a certain filling factor, with only the outflow velocities and the spatial extent as the varying parameters, and do not account for the gas mass from the luminosity of the CL. We find that this method assuming a filling factor typically overestimates the mass and kinetic energy outflow rates (Figure 15). Since the luminosity of the CLs give us the best estimate of the coronal ionized gas mass in the galaxy, in this work, we derive the coronal ionized gas mass using the luminosity of the [Si VI] CL, following the methods from 2006agna.book.....O; 2020A&ARv..28....2V; 2020ApJ...905..166L, assuming a case B recombination.
@@
The abundance ratio is obtained from 2009ARA&A..47..481A assuming a solar abundance:
The ionization fraction of [Si VI] is assumed to be 0.5: . This is taken as a conservative upper limit and we find that for lower , the change in the energetics is comparable to the errors listed in Table 4.
is determined using the line emission code PyNeb (2015A&A...573A..42L) using the CHIANTI atomic database (1997A&AS..125..149D; 2021ApJ...909...38D) using an electron temperature of 20000 K (which is the average temperatures of the gas traced by CL (2021ApJ...920...62N)) and electron density of 100 cm-3:
The calculations of the mass, momentum, and kinetic energy outflow rates depend on the above derived ionized gas mass and the spatial extent of the outflows and are given in Section 5.4.

### v2

@@ V.4 Outflows traced by Coronal Lines
Assuming the secondary component seen in the CLs is an outflow, we can derive standard estimates of energetics, including the mass outflow and kinetic energy rates. Previous studies on CL energetics (Müller-Sánchez et al., 2011; Rodríguez-Ardila et al., 2017) account for the velocities and the extents of the outflow. However, they do not measure the mass from the luminosities of the CL, but instead assume a certain filling factor and make use of the hydrogen mass alone. The unknown filling factor leads to large uncertainties in the calculated energetics. In this work, we make use of [Si VI] luminosities in order to estimate the amount of gas mass present in the outflow, and use that to determine the outflow energetics (see Appendix A for a more detailed derivation). We use the following equation to derive the outflow mass for a steady state-outflow:
@@
where is the extinction-corrected [Si VI] luminosity and is the electron density, measured from the [S II] 6716, 6731 line ratios, obtained from Gemini/GMOS spectra or Keck/LRIS spectra (Liu et al., 2020). We caution that the we adopt from the [S II] doublet likely traces lower-ionization NLR gas rather than the coronal line–emitting region itself (Baron and Netzer, 2019; Davies et al., 2020). If the CLR is denser than the [S II]-emitting gas, our assumed is underestimated and the inferred outflow masses and rates should be treated as upper limits. Additionally, the value of the outflow mass will increase proportionally if we assume sub-solar metallicities for .
@@
Note. — Column (1): Short name of the target. Column (2): Individual outflow components from the best fits. Column (3) Electron density measured from the [S II] 6716, 6731 line ratio based on the total line flux from the spatially integrated, Gemini/GMOS spectra or Keck/LRIS spectra (see Liu et al. 2020). Column (4): Outflow radius adopted in the calculation of mass, momentum, and kinetic energy outflow rates when the outflows are spatially resolved. Column (5): Average offset velocity calculated from the best fits. Column (6): Average W80 calculated from the best fits. Column (7): Ionized gas mass of the corresponding outflow component. Column (8): Ionized gas mass outflow rate of the corresponding outflow component. Column (9): Ionized gas kinetic energy outflow rate of the corresponding outflow component. Column (10): Ionized gas momentum outflow rate of the corresponding velocity component
The calculations of the mass, momentum, and kinetic energy outflow rates depend on the spatial extent of the outflows which we calculate in Section III.4 and discuss in Section V.2. Specifically, the energetics are calculated by summing up quantities over individual spaxels:
@@
In Figure 15, we plot the mass and kinetic energy outflow rates estimated for the outflowing gas traced by the CLs as a function of mass and AGN luminosity. We also compare the outflow rates from CL to the ionized [O III] gas measurements and find that the CL outflow rates are 1 dex lower, which could be due to underestimating the W80 values due to the de-wiggling process (see Section III.3). We also compare the energetics measured in this work to energetics calculated using [Si VI] outflows from ground-based low-z Seyferts (Müller-Sánchez et al., 2011; Rodríguez-Ardila et al., 2017; May et al., 2018). Since the mass outflow rates in these works were calculated assuming a filling factor, we also calculated the energetics from the AGN dwarfs using the equations presented in Müller-Sánchez et al. 2011 (indicated by the boxes with a cross). We assume a filling factor (0.001) similar to Müller-Sánchez et al. 2011 and find that this measurement over estimates the values for energetics by at least two orders of magnitude. Even if we assume higher ne and filling factor values suitable for dwarf galaxies, the energetics are still found to be higher. This could indicate that the filling factors for CL are much lower than previously assumed.
Because the highly ionized phase of the gas traced by CLs appears neither fast nor radially extended, it likely cannot expel gas from the galaxy. However, its energetics may be sufficient to disturb gas in the central regions, potentially contributing to feedback-driven gas redistribution that has been invoked to produce cored (rather than cuspy) dark-matter profiles in dwarf galaxies (de Blok, 2010). They might also accelerate gas and redistribute mass in the central, high–gravitational-potential regions, potentially driving more extended, lower-ionization outflows (e.g., [O III]).
VI Summary and Conclusions
Using JWST/NIRSpec, we undertake the first spatially resolved examination of NIR coronal lines in dwarf galaxies with AGN and ionized-gas outflows to address the challenges in unambiguously identifying AGN activity and assessing AGN-driven feedback in these low-mass systems. We generate spatial maps of the flux and velocity of the different coronal lines, allowing us to examine in detail the influence of AGN feedback, as traced by the coronal lines, on the host dwarf galaxies. Our selected targets host [O III]-traced ionized outflows and have prior evidence for CL emission, and the present study therefore focuses on a currently limited subset of dwarf galaxies. However, growing evidence for outflows in AGN dwarf galaxies (Salehirad et al., 2025) suggests that broader, more representative extensions of this work will be feasible in the future. We present our main results as follows:
@@ VI Summary and Conclusions
Based on shock and AGN models, we find that AGN photoionization is largely responsible for producing CLs, although shock excitation may contribute to the large extents of some CLs, especially if we assume lower metallicities. J1009, in particular, is dominated by AGN photoionization with minimal to no contribution from shock ionization to producing the CLs. The models also suggest that J1009 is hosting a highly accreting, low-mass black hole that is capable of producing numerous CL with I.P.s as high as 447 eV.
@@
The average extents of the CLs across the four dwarf galaxies are 0.5 kpc, which is similar to the extents found tracing NIR CL in massive galaxies. The small sizes of the dwarf galaxies in this sample mean their CLRs can occupy 10% of the host galaxy, about ten times the fractional extent seen in massive galaxies. In J0954, an outflow radius of 0.5 kpc estimated from blue-shifted UV absorption (Liu et al., 2024) agrees with the CL extents, supporting CLs as tracers of the ionized outflow.
@@
The presence, extent, and luminosity of the CLs correlate with the properties of the ionized gas outflows traced by [O III]. In galaxies with spatially resolved CL emission, CLs exhibit biconical structure, align with the outflow axis, consistent with an AGN-powered outflow. The CL luminosities correlate with both the ionized gas outflow luminosities and the FWHM of the outflows. Two of the galaxies with CL detections also have small dust extinction values, suggesting that AGN-driven outflows in these galaxies may enhance CL emission by clearing dust from the nuclear regions.
@@
Several CLs exhibit wings, suggestive of a possible secondary kinematic component. If this is an outflow, energy estimates imply this highly ionized gas in these galaxies can have a dynamical impact on the gas near galaxy centers.
@@
The results confirm that the outflows detected in the dwarf galaxies in this sample are predominantly AGN-powered and that CLs can trace their presence and impact. By revealing signatures of dust clearing, gas heating, and kinematically distinct ionized outflows, CLs may provide a direct view of how AGN-driven outflows in dwarf galaxies can exert feedback, potentially influencing their evolution. Our findings highlight the potential for CLs as diagnostics of AGN activity in low-mass, dust-poor environments, which are conditions reminiscent of galaxies in the early universe. With increased observations of AGN in faint galaxies at high redshifts (Brinchmann, 2023; Furtak et al., 2023; Harikane et al., 2023; Kocevski et al., 2023; Labbé et al., 2023; Maiolino et al., 2024), including detections of CLs such as [Ne V] in dwarf galaxies at the epoch of reionization (Chisholm et al., 2024) using JWST, understanding the formation and detection of CLs in dwarf galaxies is critical. Given that local dwarf galaxies are crucial analogs for understanding the formation and evolution of the earliest galaxies in the high-redshift universe, where they are increasingly recognised as key drivers of reionization (Atek et al., 2024) and where AGN feedback in dwarf galaxies is gaining importance in galaxy quenching (Arjona-Gálvez et al., 2024), these results can provide invaluable empirical constraints for cosmological simulations and future JWST observations aimed at unveiling the role of AGN in early galaxy evolution.
@@ Appendix A Outflow energetics
Spatially resolved IFU data has proven to be integral to deriving various energetics of outflows (the amount of mass carried by the outflows, the rates of mass, momentum and kinetic energy of the outflow), by providing constraints on the spatial extent of the outflow, which is a crucial factor in the calculation of the energetics. The energetics are dependent on several assumptions, such as the shape of the outflow (spherical vs. biconical). The ionized gas mass of the outflows can be calculated based on either the [O III] 4959, 5007 line luminosity or the Balmer line (H or H) luminosity of the outflowing, line emitting gas. Calculating the coronal gas mass using CLs is challenging. Previous works (Müller-Sánchez et al., 2011; Rodríguez-Ardila et al., 2017; May et al., 2018; Rodríguez-Ardila and Fonseca-Faria, 2020), typically assume a spherically symmetric outflow with a certain filling factor, with only the outflow velocities and the spatial extent as the varying parameters, and do not account for the gas mass from the luminosity of the CL. We find that this method assuming a filling factor typically overestimates the mass and kinetic energy outflow rates (Figure 15). Since the luminosity of the CLs give us the best estimate of the coronal ionized gas mass in the galaxy, in this work, we derive the coronal ionized gas mass using the luminosity of the [Si VI] CL, following the methods from Osterbrock and Ferland 2006; Veilleux et al. 2020; Liu et al. 2020, assuming a case B recombination.
@@
The abundance ratio is obtained from the Asplund et al. 2009 definition, given as log (Si) = log10(NSi/NH) + 12, where (NSi/NH) is the abundance ratio . For solar abundances, . If we assume a scaled-solar composition, where Si/H scales linearly with metallicity Z/Z⊙, then the abundance ratio will also scale linearly, .
The ionization fraction of [Si VI], is assumed to be 0.5. This is taken as a conservative upper limit and we find that for lower , the change in the energetics is comparable to the errors listed in Table 4.
is determined using the line emission code PyNeb (Luridiana et al., 2015) using the CHIANTI atomic database (Dere et al., 1997; Del Zanna et al., 2021) using an electron temperature of 20000 K (which is the average temperatures of the gas traced by CL (Negus et al., 2021)) and electron density of 100 cm-3:
The calculations of the mass, momentum, and kinetic energy outflow rates depend on the above derived ionized gas mass and the spatial extent of the outflows and are given in Section V.4.